# Oracle-based Uniform Sampling from Convex Bodies

### v1

@@ Appendix A Technical lemmas and proofs
Therefore, we can set $k\geq k^{*}$ where $k^{*}$ is defined in (5). The calculation for $\chi^{2}$ divergence is along the same line with the use of the second part of (4) in Theorem 2.2. This completes the proof.
The following lemma presents basic properties of Algorithm 3 and 4. We verify the unbiasedness of the rejection sampling and also provide formulas for the number of rejections.
Lemma A.3
The sample generated by Algorithm 3 and the sample generated by Algorithm 4 both follow the distribution $\pi^{X|Y}={\cal N}(y,\eta I_{d})|_{K}$.
In particular, we have that
  • For given $y\in\mathbb{R}^{d}$, let $S$ denote the event that (10) happens. Then, the acceptance rate in Algorithm 3 is
    $p_{y}=\mathrm{Pr}(S)=\frac{\int_{K}\exp\left(-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}{\int_{\mathbb{R}^{d}}\exp(-\mathcal{P}_{1}(x))\mathrm{d}x},$ (20)
    where $\mathcal{P}_{1}$ is as in (13), and the number of rejections is
    $n_{y}=\frac{1}{p_{y}}=\frac{\int_{\mathbb{R}^{d}}\exp(-\mathcal{P}_{1}(x))\mathrm{d}x}{\int_{K}\exp\left(-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}.$ (21)
  • For given $y\in\mathbb{R}^{d}$, let $T$ be the event that (16) happens. Then the acceptance rate in Algorithm 4 is
    $\mathrm{Pr}(T)=\frac{\int_{K}\exp\left(-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}{\int_{\mathbb{R}^{d}}\exp\left(-\mathcal{P}_{2}(x)\right)\mathrm{d}x}.$ (22)
    where $\mathcal{P}_{2}$ is as in (17), and the number of rejections is
    $n_{y}=\frac{\int_{\mathbb{R}^{d}}\exp\left(-\mathcal{P}_{2}(x)\right)\mathrm{d}x}{\int_{K}\exp\left(-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}.$ (23)
Proof: Let $k(x|S)$ denote the conditional density of $X$ given $S$ and $q(x)=\frac{\exp(-\mathcal{P}_{1}(x))}{\int_{\mathbb{R}^{d}}\exp(-\mathcal{P}_{1}(x))\mathrm{d}x}$ denote the density of $\mathcal{N}(\operatorname{proj}_{K}(y),\eta I_{d})$. By Bayes’ rule and (12), we have
[…]
Thus, we obtain
$\mathrm{Pr}(S)=\int\mathrm{Pr}(S|X=x)q(x)\mathrm{d}x=\frac{\int_{\mathbb{R}^{d}}\exp\left(-I_{K}(x)-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}{\int_{\mathbb{R}^{d}}\exp(-\mathcal{P}_{1}(x))\mathrm{d}x},$
which yields (20). It further leads to
$k(x|S)=\frac{\mathrm{Pr}(S|X=x)q(x)}{\mathrm{Pr}(S)}=\frac{\exp\left(-I_{K}(x)-\frac{1}{2\eta}\|x-y\|^{2}\right)}{\int_{\mathbb{R}^{d}}\exp\left(-I_{K}(x)-\frac{1}{2\eta}\|x-y\|^{2}\right)\mathrm{d}x}=\pi^{X|Y}(x|y),$
where $\pi^{X|Y}(x|y)$ is the truncated Gaussian ${\cal N}(y,\eta I_{d})|_{K}$. Therefore, we have verified $X\sim\pi^{X|Y}$. Finally, (21) holds in view of $\mathbb{E}[Z]=1/p_{y}$ for $Z\sim\text{Geom}(p_{y})$ (geometric distribution).
Verifying (22), (23) and the fact that the sample generated by Algorithm 4 follows the distribution $\pi^{X|Y}$ are similar to the above arguments and hence omitted.
Appendix B Supplementary materials for Section 4
B.1 Results about the Cutting Plane method by [27]
We first restate [27, Theorem C.1], which is about the iteration complexity and running time of the Cutting Plane method by [27].
Theorem B.1
([27, Theorem C.1]) Let $f$ be a convex function on $\mathbb{R}^{d}$. $K$ is a convex set that contains a minimizer of $f$ and $K\subseteq B_{\infty}(0,R)$, where $B_{\infty}(0,R)$ denotes a ball of radius $R$ in $\ell_{\infty}$ norm, i.e., $\left\lVert x\right\rVert_{\infty}=\sup_{1\leq i\leq d}\left|x_{i}\right\rvert$.
@@ B.1 Results about the Cutting Plane method by [27]
where $\gamma$ is as in (18).
@@
and $\gamma$ as in (18). Moreover, assume there is a separation oracle for $K$. Then the Cutting Plane method by [27] makes ${\cal O}\left(d\log\frac{d\gamma}{\alpha}\right)$ separation oracle calls to generate a $(1/d)$-solution $\hat{x}\in K$ to the optimization problem $\min_{x\in K}\frac{1}{2\eta}\left\lVert x-y\right\rVert^{2}$. In addition, we provide the following bound on $\alpha$ with high probability
@@
Proof: Since $K$ is closed per condition (A1), $K$ contains a minimizer of $f$. Moreover, the fact that $K$ is contained the Euclidean ball $B(0,R)$ per condition (A1) implies $K$ is also contain in the ball $B_{\infty}(0,R)$. Then to be able to apply Theorem B.1, we need to verify that
@@
Then Theorem B.1 guarantees that the Cutting Plane method by [27] produces a $(1/d)$-solution with $\mathcal{O}\left(d\log\frac{d\gamma}{\alpha}\right)$ separation oracle calls.
@@
Regarding the concentration inequality, recall that $y$ is the output of step 1 in the ASF (Algorithm 2) and satisfies $y=y_{k}=x_{k-1}+\sqrt{\eta}Z$, where $Z\sim{\cal N}(0,I)$ and $k$ denotes some iterate of ASF. Then, we can write
@@
The last inequality is due to $B(0,R)\supset K$ and $\operatorname{proj}_{K}(y),x_{k-1}\in K$. Combining with the Gaussian concentration inequality from [41, Equation (3.5)] to get
@@
B.2 Proof of Lemma 4.1
In view of Lemma 3.2, the fact that $\Theta^{\eta,K}_{y}(x)\geq\mathcal{P}_{1}(x)$ for every $x\in\mathbb{R}^{d}$ immediately holds, so what remains is to show $\mathcal{P}_{1}(x)\geq\mathcal{P}_{2}(x)$.
@@
This inequality and the triangle inequality imply that
@@
B.3 Proof of Proposition 4.2
The upcoming argument is similar to the proof of Proposition 3.4. Denote $\mu_{\eta}$ the distribution of $y$ at any iteration of the Algorithm 2. The average number of rejections is $\mathbb{E}_{\mu_{\eta}}[n_{y}]$ where $n_{y}$ is defined in (21). The fact that $d\mu_{\eta}/d\pi^{Y}\leq M$ from (8) implies
$\mathbb{E}_{\mu_{\eta}}[n_{y}]\leq M\mathbb{E}_{\pi^{Y}}[n_{y}],$ (33)
@@
The definition of $\mathcal{P}_{3}$} in (34) and Part b of Lemma A.1 imply
@@
where we use the fact that $2\sqrt{2/d}\leq 1/2+4/d$ in the last inequality. Consequently, applying (33) gives us the desired bound on the average number of rejections, i.e.,
$\mathbb{E}_{\mu_{\eta}}[n_{y}]\leq\sqrt{2\pi}M\exp\left(\frac{13}{4}+\frac{20}{d}\right)+M\exp\left(\frac{9}{4}+\frac{12}{d}\right).$
@@
where $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ are respectively defined at (17) and (34). The proof follows the argument showing $\mathcal{P}_{1}\geq\mathcal{P}_{2}$ in Lemma 4.1. Similar to (31), using the triangle inequality and (29), we get
@@ B.4 Further explanation for Algorithm 5
2. Generate $r\propto r^{d-1}\exp\left(-\frac{(r-b)^{2}}{2\eta}\right)$ by Adaptive Rejection Sampling for one-dimensional log-concave distribution in [20].
@@
Notice the first marginal of $p$ is $p_{r}(r)\propto r^{d-1}\exp\left(-\frac{(r-b)^{2}}{2\eta}\right)$. Due to the fact that $\log p_{r}(r)=(d-1)\log r-\frac{(r-b)^{2}}{2\eta}+\mbox{const}$ and $\frac{d^{2}}{dr^{2}}\log p_{r}(r)=-\frac{d-1}{r^{2}}-\frac{1}{\eta}<0$, $p_{r}$ is a one-dimensional log-concave density.
@@
Per the previous paragraphs, one can use any standard one-dimensional log-concave sampler (for instance [20]) to sample $r\sim p_{r}$. Then one performs uniform sampling on the $d$-dimensional unit sphere by sampling $W\sim{\cal N}(0,I)$ and set $\theta=W/\|W\|$ [55]. Finally, one outputs $X=\hat{x}+r\theta$ as the sample for $p(r,\theta)=\rho(x)$.
Appendix C Algorithms for uniform sampling on convex bodies
  • The Hit-and-Run walk is first introduced in [57] and rigorously investigated in Lovász and Simonovits in [53]. Also assuming a membership oracle on $K$ and in the special case of uniform sampling, the Hit-and-Run walk is: choose a uniform direction over the unit sphere and find a line segment in that direction that intersects $K$ at two endpoints but still belong to $K$; then go to a uniform random point on that line segment. [51] shows the its iteration complexity in total variance is of the order $\mathcal{O}\left(d^{2}C_{\mathrm{LSI}}\frac{M^{2}}{\epsilon^{2}}\right)$.
  • The In-and-Out algorithm by [39] is the ASF/proximal sampler by [42], where the RGO implentation is via a membership oracle: sample $x$ from a Gaussian distribution until $x\in K$ up to a certain number of maximum attempts $N$, at which point the algorithm halts and declares failure. Their iteration complexities in Rényi divergence and $\chi^{2}$-divergence are the same as those in our Corollary 2.3 in Section 2, except that they derive their iteration complexities via the PI constant of the uniform distribution while we use both PI and LSI constants.

### v2

@@ Appendix B Technical results and proofs
Therefore, we can take $k=O\left(\frac{1}{\eta}{C_{\mathrm{LSI}}q\log\left(2\frac{\log M}{\epsilon}\right)}\right)$ as in Theorem 2.2(a). The calculation for $\chi^{2}$ divergence is along the same line with the use of the second part of (24) in Theorem B.6. This completes the proof.
Appendix C Supplementary materials for Section 4
C.1 Results about the Cutting Plane method by [33]
We first restate [33, Theorem C.1], which is about the iteration complexity and running time of the Cutting Plane method by [33].
Theorem C.1
([33, Theorem C.1]) Let $f$ be a convex function on $\mathbb{R}^{d}$. $K$ is a convex set that contains a minimizer of $f$ and $K\subseteq B_{\infty}(0,R)$, where $B_{\infty}(0,R)$ denotes a ball of radius $R$ in $\ell_{\infty}$ norm, i.e., $\left\lVert x\right\rVert_{\infty}=\sup_{1\leq i\leq d}\left|x_{i}\right\rvert$.
@@ C.1 Results about the Cutting Plane method by [33]
where $\gamma=\frac{R}{\mbox{minwidth}(K)}$.
@@
and $\gamma=\frac{R}{\mbox{minwidth}(K)}$. Moreover, assume there is a separation oracle for $K$. Then the Cutting Plane method by [33] makes ${\cal O}\left(d\log\frac{d\gamma}{\alpha}\right)$ separation oracle calls to generate a $(1/d)$-solution $\hat{x}\in K$ to the optimization problem $\min_{x\in K}\frac{1}{2\eta}\left\lVert x-y\right\rVert^{2}$. In addition, we provide the following bound on $\alpha$ with high probability
@@
Proof: Since $K$ is closed per condition (A1), $K$ contains a minimizer of $f$. Moreover, the fact that $K$ is contained the Euclidean ball $B(0,R)$ per condition (A1) implies $K$ is also contain in the ball $B_{\infty}(0,R)$. Then to be able to apply Theorem C.1, we need to verify that
@@
Then Theorem C.1 guarantees that the Cutting Plane method by [33] produces a $(1/d)$-solution with $\mathcal{O}\left(d\log\frac{d\gamma}{\alpha}\right)$ separation oracle calls.
@@
Regarding the concentration inequality, recall that $y$ is the output of step 1 in the proximal sampler (Algorithm 2) and satisfies $y=y_{k}=x_{k-1}+\sqrt{\eta}Z$, where $Z\sim{\cal N}(0,I)$ and $k$ denotes some iterate of the proximal sampler. Then, we can write
@@
The last inequality is due to $B(0,R)\supset K$ and $\operatorname{proj}_{K}(y),x_{k-1}\in K$. Combining with the Gaussian concentration inequality from [47, Equation (3.5)] to get
@@
C.2 Proof of Lemma 4.1
In view of Lemma 3.2, the fact that $\Theta^{\eta,K}_{y}(x)\geq\mathcal{P}_{1}(x)$ for every $x\in\mathbb{R}^{d}$ immediately holds under condition (A1), so what remains is to show $\mathcal{P}_{1}(x)\geq\mathcal{P}_{2}(x)$.
@@
This inequality and the triangle inequality imply that for any generic $x\in\mathbb{R}^{d}$,
@@
C.3 Proof of Theorem 4.2
Denote $\mu^{k}$ the distribution of $y=y_{k}$ at the k-th iteration of Algorithm 2. Write $n_{y}$ the average number of proposals conditioned on $y$ in the rejection sampler employed in Algorithm 3. Then the average number of proposals is $\mathbb{E}_{\mu^{k}}[n_{y}]$ where the formula of $n_{y}$ is given in (16). The fact that $d\mu^{k}/d\pi^{Y}\leq M$ per Lemma B.4 implies
$\mathbb{E}_{\mu^{k}}[n_{y}]\leq M\mathbb{E}_{\pi^{Y}}[n_{y}],$ (34)
@@
The definition of $\mathcal{P}_{3}$} in (35) and Part b of Lemma B.1 imply
@@
where we use the fact that $2\sqrt{2/d}\leq 1/2+4/d$ in the last inequality. Consequently, applying (34) gives us the desired bound on the average number of proposals, i.e.,
$\mathbb{E}_{\mu^{k}}[n_{y}]\leq\sqrt{2\pi}M\exp\left(\frac{13}{4}+\frac{20}{d}\right)+M\exp\left(\frac{9}{4}+\frac{12}{d}\right).$
@@
where $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ are respectively defined at (13) and (35). The proof follows the argument showing $\mathcal{P}_{1}\geq\mathcal{P}_{2}$ in Lemma 4.1. Plugging in $x=y$ for (31), we obtain
Next, by the triangle inequality and the estimate (30),
[…]
which further implies
$\left\lVert y-\operatorname{proj}_{K}(y)\right\rVert^{2}$ $\leq\left\lVert y-\hat{x}\right\rVert^{2}+2\left\lVert y-\hat{x}\right\rVert\sqrt{\frac{2\eta}{d}}+\frac{2\eta}{d}$
[…]
After rearranging, this becomes
@@ C.4 A sampling subroutine
2. Generate $r\propto r^{d-1}\exp\left(-\frac{(r-b)^{2}}{2\eta}\right)$ by Adaptive Rejection Sampling for one-dimensional log-concave distribution by [26].
@@
Notice the first marginal of $p$ is $p_{r}(r)\propto r^{d-1}\exp\left(-\frac{(r-b)^{2}}{2\eta}\right)$. Due to the fact that $\log p_{r}(r)=(d-1)\log r-\frac{(r-b)^{2}}{2\eta}+\mbox{const}$ and $\frac{\mathrm{d}^{2}}{\mathrm{d}r^{2}}\log p_{r}(r)=-\frac{d-1}{r^{2}}-\frac{1}{\eta}<0$, $p_{r}$ is a one-dimensional log-concave density.
@@
Per the previous paragraphs, one can use any standard one-dimensional log-concave sampler, for instance [26], to sample $r\sim p_{r}$. Then one performs uniform sampling on the $d$-dimensional unit sphere by sampling $W\sim{\cal N}(0,I)$ and setting $\theta=W/\|W\|$ [63]. Finally, one outputs $X=\hat{x}+r\theta$ as the sample for $p(r,\theta)=\rho(x)$.
Appendix D Examples of convex bodies and their oracle implementations
This section provides details on three types of convex bodies, namely general Z-polytopes, general H-polytopes, and bounded spectrahedron, and their corresponding oracle implementations (i.e., membership, separation, and projection). In particular, the three oracles have the same arithmetic complexity for general Z-polytopes and bounded spectrahedron.
General Z-Polytopes.
A Z-polytope in $\mathbb{R}^{d}$ is a polytope that can be written as a Minkowski sum of finitely many line segments, or equivalently as a linear image of a hypercube:
$Z=\left\{\sum_{j=1}^{m}t_{j}v_{j}:t_{j}\in[-1,1]\right\}=V[-1,1]^{m},\qquad V=[v_{1}\dots v_{m}]\in\mathbb{R}^{d\times m},$
see, e.g., [66].
When $m=d$ and $V$ is invertible, $Z=V[-1,1]^{d}$ is a parallelepiped, and the map $U\sim\mathrm{Unif}([-1,1]^{d})$, $X:=VU$ is a bijection with constant Jacobian determinant, so $X$ is uniform on $Z$. Thus parallelepipeds are as easy to sample from as hypercubes. For general Z-polytopes, $t\mapsto Vt$ is not injective and uniform $t$ on $1,1]^{m}$ does not induce uniform measure on $Z$, so uniform sampling is non-trivial and typically done via random-walk or MCMC methods e.g., [7].
For general Z-polytopes, membership, separation, and projection oracles can all be implemented with comparable arithmetic complexity $\mathcal{O}((m+d)^{3})$ using interior-point methods.
  • A membership oracle for $Z$ takes $x\in\mathbb{R}^{d}$ and decides whether $x\in Z$. Membership reduces to the box-constrained feasibility problem: find $t\in\mathbb{R}^{m}$ such that $Vt=x$ and $\mathbf{1}\leq t\leq\mathbf{1}$. This can be formulated as a linear program (LP) and solved by interior-point methods. Generic primal-dual interior-point methods for dense LPs have arithmetic complexity $\mathcal{O}((m+d)^{3})$; see [65].
  • A separation oracle for $Z$ takes $x\in\mathbb{R}^{d}$ and, if $x\notin Z$, returns a halfspace $\{y:a^{\top}y\leq b\}$ containing $Z$ but not $x$. If the LP above is infeasible, its dual (via Farkas’ lemma) yields such a separating hyperplane (see, e.g., [3, Exercise 5.23]). Thus a separation oracle can be implemented with essentially the same complexity as the LP-based membership oracle: one primal-dual LP solve, again $\mathcal{O}((m+d)^{3})$.
  • A projection oracle for $Z$ maps $x\in\mathbb{R}^{d}$ to $\Pi_{Z}(x):=\arg\min_{z\in Z}\|z-x\|_{2}$. Hence projection reduces to a convex quadratic program (QP) with box constraints. Such QPs admit polynomial-time interior-point methods with arithmetic complexity $\mathcal{O}((m+d)^{3})$ for generic dense instances [65].
General H-Polytopes.
An $H$-polytope in $\mathbb{R}^{d}$ has the form
$P=\{x\in\mathbb{R}^{d}:Ax\leq b\},$
where $A\in\mathbb{R}^{m\times d}$, $b\in\mathbb{R}^{m}$, with componentwise inequalities.
This class includes some simple convex bodies with explicit uniform samplers: (a) the hypercube $1,1]^{d}$, sampled by drawing coordinates independently from $\mathrm{Unif}[-1,1]$; (b) the standard simplex $\Delta^{d-1}=\{x\in\mathbb{R}^{d}:x_{i}\geq 0,\ \sum_{i}x_{i}=1\}$, sampled via independent exponentials followed by normalization; see [18, Chapter 9]. These examples have special product or radial structure. For a general H-polytope with no special structure, uniform sampling usually relies on MCMC (ball walk, hit-and-run, Dikin walk, etc.; see [22, 60]).
For general H-polytopes, membership and separation oracles have complexity $\mathcal{O}(md)$, while projection is costlier.
  • A membership oracle takes $x\in\mathbb{R}^{d}$ and checks $Ax\leq b$. Computing $Ax$ costs $\mathcal{O}(md)$ and comparisons cost $\mathcal{O}(m)$, so membership runs in $\mathcal{O}(md)$ time.
  • A separation oracle for $P$ takes $x\in\mathbb{R}^{d}$ and either certifies $x\in P$ or returns a halfspace containing $P$ but not $x$. If $x\notin P$, some inequality is violated, say $(Ax)_{i}>b_{i}$. Let $a:=A_{i,\cdot}$ and $\beta:=b_{i}$. Then every $y\in P$ satisfies $a^{\top}y\leq\beta$ while $a^{\top}x>\beta$, so the hyperplane $\{y:a^{\top}y=\beta\}$ separates $x$ from $P$. Thus separation costs $\mathcal{O}(md)$, the same order as membership.
  • A projection oracle for $P$ maps $x\in\mathbb{R}^{d}$ to $\Pi_{P}(x):=\arg\min_{z\in\mathbb{R}^{d}}\tfrac{1}{2}\|z-x\|_{2}^{2}$ s.t. $Az\leq b$. This is a convex QP with linear inequality constraints. Generic polynomial-time algorithms via interior-point methods have worst-case arithmetic complexity $\mathcal{O}((m+d)^{3})$ for dense problems [65].
Bounded Spectrahedron.
Let $\mathbb{S}^{n}$ be the space of real symmetric $n\times n$ matrices with Frobenius inner product $\langle A,B\rangle=\mathrm{Tr}(A^{\top}B)$. The positive semidefinite cone $\mathbb{S}^{n}_{+}:=\{X\in\mathbb{S}^{n}:X\succeq 0\}$ is closed and convex [3, Section 2.2.4].
For a fixed radius $R>0$, consider the bounded spectrahedron
$K_{R}:=\{X\in\mathbb{S}^{n}:X\succeq 0,\ \mathrm{Tr}(X)\leq R\}.$
It is the intersection of $\mathbb{S}^{n}_{+}$ with an affine trace constraint, hence convex. For $X\succeq 0$ we have $\|X\|_{F}^{2}=\sum_{i}\lambda_{i}^{2}\leq(\sum_{i}\lambda_{i})^{2}=\mathrm{Tr}(X)^{2}\leq R^{2}$, so $K_{R}$ is bounded. Being also closed, $K_{R}$ is compact. Via the identification of $\mathbb{S}^{n}$ with $\mathbb{R}^{d}$, $d=n(n+1)/2$, equipped with $\|\cdot\|_{F}$, $K_{R}$ is a compact convex body in a Euclidean space.
Membership, separation, and projection oracles on $K_{R}\subset\mathbb{S}^{n}$ can all be implemented in time $\mathcal{O}(n^{3})$, dominated by eigenvalue decompositions.
  • Membership check is performed as followed: given $X\in\mathbb{S}^{n}$, compute an eigendecomposition $X=Q\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})Q^{\top}$ and check $\lambda_{i}\geq 0$ and $\sum_{i=1}^{n}\lambda_{i}\leq R$. The sum is $\mathcal{O}(n)$, while symmetric eigendecomposition for dense $X$ costs $\mathcal{O}(n^{3})$ [70], so membership is $\mathcal{O}(n^{3})$.
  • Regarding the separation oracle, if $X\notin K_{R}$, either $X$ is not PSD (some $\lambda_{\min}<0$) or $X\succeq 0$ but $\mathrm{Tr}(X)>R$.
    PSD violation: Let $(\lambda_{\min},v)$ be an eigenpair with $\lambda_{\min}<0$ and $\|v\|_{2}=1$, set $N:=vv^{\top}$, and define $\varphi(Y):=\langle N,Y\rangle=v^{\top}Yv$. For any $Y\succeq 0$ we have $v^{\top}Yv\geq 0$, so every $Y\in K_{R}$ satisfies $\langle N,Y\rangle\geq 0$, i.e., $K_{R}\subseteq\{Y:\langle N,Y\rangle\geq 0\}$. For $X$ we have $\langle N,X\rangle=v^{\top}Xv=\lambda_{\min}<0$, so the hyperplane $\{Y:\langle N,Y\rangle=0\}$ strictly separates $X$ from $K_{R}$.
    Trace violation: If $X\succeq 0$ but $\mathrm{Tr}(X)>R$, take $N:=I_{n}$ and $\psi(Y):=\langle N,Y\rangle=\mathrm{Tr}(Y)$. Every $Y\in K_{R}$ satisfies $\mathrm{Tr}(Y)\leq R$, i.e. $\langle N,Y\rangle\leq R$, so $K_{R}\subseteq\{Y:\langle N,Y\rangle\leq R\}$. For $X$ we have $\langle N,X\rangle=\mathrm{Tr}(X)>R$, so the hyperplane $\{Y:\langle N,Y\rangle=R\}$ strictly separates $X$ from $K_{R}$.
    In both cases, the separator uses eigeninformation already computed for membership, so separation also costs $\mathcal{O}(n^{3})$.
[…]
    Thus, a projection oracle for $K_{R}$ is computed as follows: 1) compute $V=Q\mathrm{diag}(\lambda)Q^{\top}$ (cost $\mathcal{O}(n^{3})$); 2) project $\lambda$ onto $C_{R}$ to obtain $\mu$; and 3) return $\Pi_{K_{R}}(V)=Q\mathrm{diag}(\mu)Q^{\top}$. The overall complexity dominated by the eigendecomposition is $\mathcal{O}(n^{3})$.